Integration Procedures(Itevrations by Parts)


∫ x³ cos 5x dx


1
Expert's answer
2021-04-15T07:27:13-0400

The basic formula is


"\displaystyle\int u\ dv=uv-\int v\ du"


The formula may be applied again if necessary


"\displaystyle\int x^3 \cos 5x\ dx="

denote "u=x^3" and "dv=\cos 5x\ dx"

then "du=3x^2 dx" and "v=\dfrac{1}{5}\sin5x"


"\displaystyle x^3\cdot\dfrac{1}{5}\sin5x-\int \dfrac{1}{5}\sin5x\cdot3x^2 dx="

simplify

"\displaystyle \dfrac{1}{5}x^3\sin5x-\dfrac{3}{5}\int \sin5x\cdot x^2 dx="

"\displaystyle \dfrac{1}{5}x^3\sin5x-\dfrac{3}{5}\int x^2\cdot\sin5x dx="

integate by parts again

denote "u=x^2" and "dv=\sin5x\ dx"

then "du=2x dx" and "v=-\dfrac{1}{5}\cos5x"


"\displaystyle \dfrac{1}{5}x^3\sin5x-\dfrac{3}{5}\left(x^2\cdot\left(-\dfrac{1}{5}\cos5x\right)-\int \left(-\dfrac{1}{5}\cos5x\right)\cdot2x dx\right)="


"\displaystyle \dfrac{1}{5}x^3\sin5x-\dfrac{3}{5}\left(-\dfrac{1}{5}x^2\cos5x+\dfrac{2}{5}\int \cos5x\cdot x dx\right)="


"\displaystyle \dfrac{1}{5}x^3\sin5x+\dfrac{3}{25}x^2\cos5x-\dfrac{6}{25}\int x\cdot\cos5x dx="


integrate by parts again

denote "u=x" and "dv=\cos 5x\ dx"

then "du=dx" and "v=\dfrac{1}{5}\sin5x"


"\displaystyle \dfrac{1}{5}x^3\sin5x+\dfrac{3}{25}x^2\cos5x-\dfrac{6}{25}\left(x\cdot\dfrac{1}{5}\sin5x-\int \dfrac{1}{5}\sin5x\cdot dx\right)="


"\displaystyle \dfrac{1}{5}x^3\sin5x+\dfrac{3}{25}x^2\cos5x-\dfrac{6}{25}\left(\dfrac{1}{5}x\sin5x- \dfrac{1}{5}\int\sin5x\ dx\right)="


"\displaystyle \dfrac{1}{5}x^3\sin5x+\dfrac{3}{25}x^2\cos5x-\dfrac{6}{125}x\sin5x+\dfrac{6}{125}\int\sin5x\ dx="


"\displaystyle \dfrac{1}{5}x^3\sin5x+\dfrac{3}{25}x^2\cos5x-\dfrac{6}{125}x\sin5x+\dfrac{6}{125}\left(-\dfrac{1}{5}\cos5x\right)+c="


"\displaystyle \dfrac{1}{5}x^3\sin5x+\dfrac{3}{25}x^2\cos5x-\dfrac{6}{125}x\sin5x-\dfrac{6}{625}\cos5x+c"